OpenStudy (anonymous):

Simplify. Express the quotient as a radical expression. ninth root of x to the seventh power over ninth root of x to the fifth power

OpenStudy (dumbcow):

\[\large \frac{\sqrt[9]{x^{7}}}{\sqrt[9]{x^{5}}} = x^{7/9-5/9} = x^{2/9} = \sqrt[9]{x^{2}}\]

Directrix (directrix):